OpenStudy (anonymous):

Solve the Inequality 1/4c<-8

OpenStudy (texaschic101):

try dividing both sides by 1/4

OpenStudy (texaschic101):

or multiplying both sides by 4

OpenStudy (perl):

$$ \Large { \frac{1}{4}c < - 8\\~\\ \color{red}{\frac 4 1} \cdot \frac{1}{4}c < - 8 \cdot \color{red}{\frac 4 1}\\ } $$